# DualLQR: Efficient Grasping of Oscillating Apples using Task Parameterized Learning from Demonstration

### v1

@@ II-B During execution
Fig. 2 shows the steps during execution. The start frame is shown in blue and the end frame, which can be oscillating, is shown in orange. The first step is to transform the task from the world frame, shown in Fig. 2a, to the two reference frames, shown in Fig. 2b. In each reference frame $i$, the end-effector pose $x_{i,t}$ was provided to the LQR, shown in Fig. 2c. Since the LQRs were fitted before execution, each immediately calculated the control output. The last step was to combine the control output of each LQR, shown in Fig. 2d. The control outputs were transformed back into the world frame and the weighted average is calculated using the covariance matrix $\Sigma_{i,t}$ from each GMR $i$. Since we aimed to prioritise the reference frame with the least variance, we used the inverse covariance matrix, i.e. the precision matrix. This resulted in the following equation:
$\begin{split}u_{t}=\left(\text{diag}\left(\Sigma_{1,t}^{-1}\right)+\text{diag}\left(\Sigma_{2,t}^{-1}\right)\right)^{-1}\cdot\\ \left(\text{diag}\left(\Sigma_{1,t}^{-1}\right)\cdot u_{1,t}+\text{diag}\left(\Sigma_{2,t}^{-1}\right)\cdot u_{2,t}\right)\end{split}$ (1)
@@
With $u_{i,t}$ being the vector containing controls of LQR $i$ at timestep $t$ and $\Sigma_{i,t}$ being the covariance matrix of GMR $i$ at timestep $t$. In our case, $i$ can be $1$ and $2$, indicating the start and reference frames. Lastly, $u_{t}$ is the vector containing the combined and weighted controls. The covariance matrix provided by the GMR is non-sparse. To ensure control output in one dimension does not influence control output in other dimensions, the covariance matrix is changed to a diagonal matrix by setting all non-diagonal values to zero.
From the combined controls and the current end-effector pose, the next end-effector pose was calculated. The last step was calculating the inverse kinematics to control the robot arm. This was done using the Levenberg-Marquardt Numerical Inverse Kinematics Solver.
III Experiments
@@
For both experiments, we used a model trained on 40 demonstrations. In these demonstrations, the target was a 10cm cube, which was static. Using a static target in the demonstrations simplified the process of collecting demonstrations. Fig. 3 shows the path of a typical demonstration, going from $t=0$ to $t=T$. The same demonstration dataset was used as in Van de Ven et al. [19].
@@ III-A Simulation experiment
In the simulation experiment, we tested the infinite horizon LQR (InfLQR) [16] and our proposed DualLQR. The simulation was created using ROS2, using a simulated UR5e manipulator. The moving target was created by defining a target pose with sinusoidal movement on one dimension. In this experiment, we adjusted the control cost, the axis along which the oscillation took place and the amplitude of the oscillation. The control cost allowed optimization between the closeness of tracking and the total distance travelled by the manipulator. Through using different axes and amplitudes, we investigated the robustness and which oscillations were the most challenging for the methods. The control cost was adjusted between $-3.0$ and $3.0$ in steps of $0.3$. The oscillation was performed along x, y, z, roll, pitch, and yaw. For the results, we grouped these into position or orientation oscillations. The oscillation amplitude was set at four levels, specifically no oscillation, low oscillation of 0.05 m or 0.15 rad, medium oscillation of 0.10 m or 0.30 rad, and high oscillation of 0.15 m or 0.45 rad.
@@
In each test, we performed eleven repetitions, placing the goal in a new pose each time. We used one central pose at $[0.22,0.27,-0.26,0.00,0.00,1.46]$. The remaining ten poses changed one dimension, using the following values: X: ±$\pm 0.1$, Y: ±$\pm 0.2$, Z: ±$\pm 0.05$, roll: ±$\pm 0.2$, yaw: ±$\pm 0.2$. These poses were based on the variation in the demonstration dataset, which had minimal variation along the pitch axis.
@@
The task of apple harvesting requires high accuracy in the final approach, combined with a short path. To calculate the accuracy in the final approach, the first step was to define the final approach. The start of the final approach is indicated in Fig. 3 using a purple line. In the end-frame, this point lies at Y=0.05 m. This point was chosen as it indicates the position from where collisions with the target can occur, which would result in failed grasps. As all demonstrations were successful, we used the median absolute deviations from the origin as the limits for the final approach phase. Staying below these limits indicates a successful final approach phase in the trajectory of an execution. These limits were 0.03 m along the Y-axis, 0.10 m along the Z-axis, 0.07 rad along the roll axis, 0.07 rad along the pitch axis, and 0.07 along the yaw axis. For each execution, we determined at what timestep the final approach started. From there, we calculated the fraction of timesteps where the end-effector pose was within these limits. We used the fraction in the demonstrations, which was at 0.88, as the required accuracy. For the requirement of a short path, we looked at the distance the manipulator travelled to perform the task. This was calculated using the Euclidean distance in meters and the quaternion distance. These were calculated for all timesteps. This results in the following equation:
$d=\sum_{t=1}^{T}\left(\left|p_{t}-p_{t+1}\right|+\min\left(\left|q_{t}-q_{t+1}\right|,\left|q_{t}+q_{t+1}\right|\right)\right)$ (2)
With $p$ being the position vector and $q$ being the orientation, represented as quaternion.
@@
III-B Apple grasping experiment
Next, we tested the performance of the method in a real-world environment while doing a harvesting task, specifically apple grasping using a suction cup gripper. Fig. 4 shows the initial state of the apple grasping set-up. On the left, the UR5e manipulator is shown. The end-effector is a suction cup. On the right, the apple is shown. We used a fake apple for the experiments, with OptiTrack [20] markers to track its pose. The apple was attached to a rod, oscillating in a pendulum motion, along the X-axis of the end-frame. Before each run, the apple was positioned at one of the extremes of the pendulum motion. At the start of each run, the apple was released, resulting in a dampened oscillating motion. Compared to the simulation experiment, the oscillation of the apple corresponds to the combination of a medium position oscillation and a medium orientation oscillation.
@@ III-B Apple grasping experiment
The goal of this experiment was to grasp the apple successfully. In addition to the aforementioned metrics of final accuracy and distance travelled, we evaluated the time until successful grasping. A successful grasp happened when the air pressure in the suction cup dropped below 0.9 bar. If this happened before the trajectory was completed, the time the pressure dropped below 0.9 bar was used as the time until successful grasping. Otherwise, the complete time to execute the trajectory was stored.
IV Results
IV-A Results simulation experiment
(a) Results of the simulation testing of the InfLQR.
(b) Results of the simulation testing of the Dual LQR.
@@ IV-A Results simulation experiment
Figure 5: Results of the simulation testing of both methods. Darker lines indicate higher oscillation levels. The levels are described in section III-A. The figures on the left show the effect of an oscillation of the target’s position and the figures on the right show the effect of an oscillation of the target’s orientation. At each tested setting, a vertical bar indicates the 95% confidence interval. The black dashed line indicates the required accuracy, set at 0.88.
Fig. 5a shows the performance of InfLQR in terms of accuracy and distance travelled as a function of the control cost used. With increasing position oscillations, the accuracy dropped slightly. With high oscillations, only two controller settings reached the required accuracy, specifically at a control cost of $0.0$ and $-3.0$. A slight effect of control cost was observed, with higher control costs resulting in lower accuracy. The InfLQR was unable to correctly deal with orientation oscillations. From low oscillations, the method was unable the reach the required accuracy for any control cost. In addition, higher control costs reduced the accuracy. Each level of oscillation added more distance travelled by the manipulator. This effect was bigger for orientation oscillations, and reduced by increasing the control cost.
@@
Fig. 5b shows the performance of the DualLQR in terms of accuracy and distance travelled as a function of the control cost used. The effect of the oscillations on the accuracy was much smaller compared to the InfLQR. Without oscillations, it can be seen that high control costs resulted in lower accuracy. To reach the required accuracy, the control cost needed to be $2.4$ or lower. Introducing oscillations along the position of the target had a minimal effect on the accuracy. The value of control cost from which the required accuracy was reached dropped slightly. For low and medium oscillations, this was $2.1$ or lower. For high oscillations, this was $1.8$ or lower. Introducing oscillations along the orientation of the target resulted in a similar behaviour, but the accuracy of low control costs also dropped slightly. For low oscillations, the requirement was met at a control cost of $2.1$ or lower. For medium oscillations, the controller reached the requirement from a control cost of $0.3$ or lower, except for $-0.3$. For high oscillations, the controller reached the requirement from a control cost of $0.0$ or lower, except for $-0.3$. Each level of oscillation added more distance travelled by the manipulator. This effect was bigger for orientation oscillations, and reduced by increasing the control cost.
@@
To compare the methods, we selected the highest control cost that reached the required accuracy. For the InfLQR, we selected a control cost of $0.0$, which was the highest control cost that reached the required accuracy under high position oscillations. For the DualLQR, we selected a control cost of $0.0$ as well, this was the highest control cost that reached the required accuracy under high orientation oscillations. Table I shows the performance of both methods for all oscillations. DualLQR resulted in higher accuracy for all oscillations, increasing on average 18%. Only medium and high position oscillations did not result in significantly higher accuracy. For the Euclidean distance, similar results were observed, with an average decrease of 3.7%. For rotated distance, the DualLQR decreased the distance with an average of 2.8%. DualLQR performed better for position oscillations, while the InfLQR was better for high oscillations along the orientation. However, the much lower accuracy indicates worse tracking, which reduced the distance.
@@
TABLE I: Performance in simulation of best models for each oscillation level & method. Bold values indicate the best method when there is a significant difference between both methods at a p-value of 0.05.
@@
None DualLQR 0.96$\pm$0.11 0.59$\pm$0.039 0.85$\pm$0.036
InfLQR 0.91$\pm$0.15 0.62$\pm$0.033 0.89$\pm$0.038
Low, Ori DualLQR 0.95$\pm$0.16 0.61$\pm$0.041 1.21$\pm$0.060
InfLQR 0.83$\pm$0.23 0.63$\pm$0.033 1.24$\pm$0.057
Low, Pos DualLQR 0.98$\pm$0.10 0.73$\pm$0.043 0.84$\pm$0.036
InfLQR 0.90$\pm$0.15 0.76$\pm$0.039 0.89$\pm$0.038
Medium, Ori DualLQR 0.92$\pm$0.15 0.66$\pm$0.081 1.71$\pm$0.131
InfLQR 0.72$\pm$0.20 0.66$\pm$0.039 1.69$\pm$0.096
Medium, Pos DualLQR 0.95$\pm$0.16 0.94$\pm$0.102 0.85$\pm$0.045
InfLQR 0.91$\pm$0.19 1.00$\pm$0.046 0.89$\pm$0.040
High, Ori DualLQR 0.88$\pm$0.13 0.65$\pm$0.044 2.22$\pm$0.140
InfLQR 0.55$\pm$0.25 0.69$\pm$0.053 2.14$\pm$0.135
High, Pos DualLQR 0.96$\pm$0.14 1.20$\pm$0.078 0.84$\pm$0.047
InfLQR 0.91$\pm$0.14 1.23$\pm$0.054 0.91$\pm$0.039
IV-B Results apple grasping experiment
When testing the performance of DualLQR on apple grasping, we found a very high success rate. Only a single grasp attempt out of 110 was unsuccessful, at a control cost of $-0.3$, shown with a cross in Fig. 6. For this control cost, the final approach accuracy was also insufficient, as shown in Fig. 6. From -0.9, the required accuracy was met, except for -2.1. At this control cost, the apple yaw of the apple was around $0.3$ less than the other settings.
@@ IV-B Results apple grasping experiment
In this experiment, the control cost needed to be lower to achieve the required accuracy, as shown in the top plot in Fig. 6. However, similar accuracy was achieved with low control costs. Furthermore, the distance travelled by the manipulator was nearly doubled compared to the simulation, as shown in the middle plot in Fig. 6. Lastly, the bottom plot of Fig. 6 shows that the time until successful grasping was the lowest at a control cost of $-1.2$. At a control cost of $-1.8$ or lower, there was a lot of variation, while control costs above $-1.2$ were slightly slower.

### v3

@@ 3.0.2 During execution
At each timestep and for each reference frame, the control output can be directly determined using Equation 17. The outputs of each reference frame are transformed to be in the global reference frame using $A_{i}$ from Equation 1. In order to combine the control outputs, the weighted average was calculated using the covariance matrix $\Sigma_{i,t}$ from each GMR $i$. Since we aimed to prioritise the reference frame with the least variance, we used the inverse covariance matrix, i.e. the precision matrix. This resulted in the following equation:
$u_{t}=\frac{\sum_{i=1}^{I}\text{diag}\left(\left(\hat{\Sigma}_{i,t}^{\mathcal{O}}\right)^{-1}\right)u_{i,t}}{\sum_{i=1}^{I}\text{diag}\left(\left(\hat{\Sigma}_{i,t}^{\mathcal{O}}\right)^{-1}\right)}$ (18)
@@
with $u_{i,t}$ being the vector containing controls of LQR $i$ at timestep $t$ and $\Sigma_{i,t}^{\mathcal{O}}$ being the covariance matrix of GMR $i$ at timestep $t$. In our case, $i$ can be $1$ and $2$, indicating the start and end reference frames. Lastly, $u_{t}$ is the vector containing the combined and weighted controls. The covariance matrix provided by the GMR is non-sparse, which is used for the fitting of the LQR. For combining the outputs, only the weighted average needs to be determined. As the control outputs are determined for the correct output dimension, the weighted average should not affect this. Therefore, the covariance matrices were changed to diagonal matrices by setting all non-diagonal values to zero. This follows the convention of matrix $B$ of Equation 8. The steps performed during execution are shown in Fig. 1b.
4 Experimental set-up
@@
For both experiments, we used a model trained on 40 demonstrations. In these demonstrations, the target was a 10 cm cube, which was static. A static target was used both to simplify the process of collecting demonstrations, and because the variable conditions in the orchard can result in static targets in all demonstrations. Fig. 2 shows the path of a typical demonstration, going from $t=0$ to $t=T$. The same demonstration dataset was used as in Van de Ven et al. [20].
@@ 4.1 Set-up of simulation experiment
In the simulation experiment, we tested the infinite horizon LQR (InfLQR) [4] and our proposed DualLQR. The simulation was created using ROS2, using a simulated UR5e manipulator. The moving target was created by defining a target pose with sinusoidal movement in one dimension. In this experiment, we adjusted the control cost $\rho$, the axis along which the oscillation took place and the amplitude of the oscillation. The control cost $\rho$ allowed optimization between the closeness of tracking and the total distance travelled by the manipulator, as indicated in Equation 9. The control cost $\rho$ was adjusted between $-3.0$ and $3.0$ in steps of $0.3$. Through using different movement axes and amplitudes, we investigated the robustness and which oscillations were the most challenging for the methods. The oscillation was performed along x, y, z, roll, pitch, and yaw in the object coordinate frame. For grasping apples, the orientation is also important. While the apple itself is roughly spherical, it is attached at the peduncle, and requires a rotation around this peduncle for correct detachment. While any orientation of the end-effector could still be used to grasp the fruit, the detachment motion has the least influence on the workspace if it can be performed with the final joint of the manipulator. Because of this, the orientation of the fruit needs to be tracked. For the results, we grouped these into position or orientation oscillations. The oscillation amplitude was set at four levels, specifically no oscillation, low oscillation of 0.05 m or 0.15 rad, medium oscillation of 0.10 m or 0.30 rad, and high oscillation of 0.15 m or 0.45 rad.
@@
In each test, we performed eleven repetitions, placing the goal in a new pose each time. We used one central pose at $[0.22,0.27,-0.26,0.00,0.00,1.46]$. The remaining ten poses changed one dimension, using the following values: X: ±$\pm 0.1$, Y: ±$\pm 0.2$, Z: ±$\pm 0.05$, roll: ±$\pm 0.2$, yaw: ±$\pm 0.2$. These poses were based on the variation in the demonstration dataset, where there was minimal variation along the pitch axis.
(a) Top view
(b) Side view
Figure 3: Example of final approach. The wireframes indicate the end-frame. The black line indicates the trajectory. The purple line at 5 cm from the target indicates the start of the final approach. The orange lines indicate the boundaries. The trajectory is dashed green if it lies within the boundaries. Else, it is dashed red.
@@
The task of apple harvesting requires high accuracy in the final approach, combined with a short path. To calculate the accuracy in the final approach, the first step was to define the final approach. The start of the final approach is indicated in Fig. 2 using a purple line. In the end-frame, this point lies at Y=0.05 m. This point was chosen as it indicates the position from where collisions with the target can occur, which would result in failed grasps. A close-up of the final approach is shown in Fig.3. Once the end-effector is closer than 5 cm along the Y-axis, the remaining section is the final approach of the trajectory. Within the final approach, the pose of each timestep is evaluated for being within the accuracy boundaries, indicated in orange. If the pose is within the boundaries for all dimensions, that pose is counted as accurate. The final approach accuracy is the number of accurate poses divided by the total number of poses in the final approach. As all demonstrations were successful, we used the median absolute deviations from the origin as the limits for the final approach phase. Staying below these limits indicates a successful final approach phase in the trajectory of an execution. These limits were 0.03 m along the X-axis, 0.10 m along the Z-axis, 0.07 rad along the roll axis, 0.07 rad along the pitch axis, and 0.07 along the yaw axis. We used the fraction in the demonstrations, which was at 0.88, as the required accuracy. For the requirement of a short path, we looked at the distance the manipulator travelled to perform the task. This was calculated using the Euclidean distance in meters and the quaternion distance in radians. These were calculated for all timesteps. This results in the following equation:
$d=\sum_{t=1}^{T}\left(|p_{t}-p_{t+1}\rvert+2\arccos\left(|q_{t}\cdot q_{t+1}\rvert\right)\right)$ (19)
With $p$ being the position vector and $q$ being the orientation, represented as a quaternion.
@@
4.2 Set-up of apple grasping experiment
Next, we tested the performance of the method in a real-world environment while doing a harvesting task, specifically apple grasping using a suction cup gripper, using the same demonstration set. Fig. 4 shows the initial state of the apple grasping set-up. On the left, the UR5e manipulator is shown. The end-effector is a suction cup. On the right, the apple is shown. We used a fake apple for the experiments, with OptiTrack [8] markers to track its pose. The apple was attached to a rod, oscillating in a pendulum motion, along the X-axis of the end-frame. Before each run, the apple was positioned at one of the extremes of the pendulum motion. At the start of each run, the apple was released, resulting in a dampened oscillating motion. Compared to the simulation experiment, the oscillation of the apple corresponds to the combination of a medium position oscillation and a medium orientation oscillation.
@@ 4.2 Set-up of apple grasping experiment
The goal of this experiment was to grasp the apple successfully. In addition to the aforementioned metrics of final accuracy and distance travelled, we evaluated the time until successful grasping. A successful grasp happened when the air pressure in the suction cup dropped below 0.9 bar. If this happened before the trajectory was completed, the time the pressure dropped was used as the time until grasping. Otherwise, the time to execute the trajectory was stored.
5 Results
5.1 Results of simulation experiment
(a) InfLQR
(b) Single LQR
(c) Dual LQR
@@ 5.1 Results of simulation experiment
Figure 5: Results of the simulation testing of both methods. Darker lines indicate higher oscillation levels. The levels are described in section 4.1. The figures on the left show the effect of an oscillation of the target’s position, and the figures on the right show the effect of an oscillation of the target’s orientation. At each tested setting, a vertical bar indicates the 95% confidence interval. The black dashed line indicates the required accuracy, set at 0.88.
Fig. 5 shows the performance of InfLQR, a single LQR and DualLQR in terms of accuracy and distance travelled as a function of the control cost $\rho$ used.
For all controllers, increasing oscillations increased the distance travelled, and increasing the control cost reduced the distance travelled. The single LQR had the largest distance increase, and InfLQR and DualLQR were quite similar.
@@
For InfLQR, with increasing position oscillations, the accuracy dropped slightly. With high oscillations, only two controller settings reached the required accuracy. A slight effect of control cost was observed, with higher control costs resulting in lower accuracy. The InfLQR was unable to reach the required accuracy for any control cost. For a single LQR, high values of the control cost resulted in reduced accuracy in the final approach. Without oscillations, the control cost needed to be $1.8$ or lower. Increased position oscillations had a minimal influence. Only with high oscillations the control cost needed to be $1.5$ or lower. Introducing oscillations around the orientation axes resulted in similar behaviour, but the effect on accuracy and distance was greater. At high oscillations, the required accuracy could not be reached. For DualLQR, the effect of the oscillations on the accuracy was the smallest. Without oscillations, high control costs resulted in lower accuracy. To reach the required accuracy, the control cost needed to be $2.4$ or lower. Introducing oscillations along the position of the target had a minimal effect. The value of control cost from which the required accuracy was reached dropped. For low and medium oscillations, this was $2.1$. For high oscillations, this was $1.8$. Introducing oscillations along the orientation of the target resulted in a similar behaviour, but also affected the accuracy of low control costs. For low oscillations, the requirement was met at a control cost of $2.1$ or lower. For medium oscillations, the controller reached the requirement from a control cost of $0.3$ or lower, except for $-0.3$. For high oscillations, the controller reached the requirement from a control cost of $0.0$ or lower, except for $-0.3$.
@@
To compare the methods, we selected the highest control cost that reached the required accuracy. For the InfLQR, we selected a control cost of $0.0$, which was the highest control cost that reached the required accuracy under high position oscillations. For the single LQR, we selected a control cost of $-0.3$, which was the control cost that reach the required accuracy under medium orientation oscillations. For the DualLQR, we selected a control cost of $0.0$ as well, this was the highest control cost that reached the required accuracy under high orientation oscillations. Table 1 shows the performance of both methods for all oscillations. DualLQR resulted in higher accuracy for all oscillations, increasing on average 18% compared to InfLQR and 7.0% compared to a single LQR. For the Euclidean distance, similar results were observed, with an average decrease of 3.9% compared to InfLQR and 26% compared to a single LQR. For rotated distance, the DualLQR decreased the distance with an average of 2.8% compared to InfLQR and 6.8% compared to a single LQR. The InfLQR had lower rotated distance for medium and high orientation oscillations, but this also resulted in much lower final approach accuracy.
@@
Table 1: Performance in simulation of best models for each oscillation level & method. Each value shows the average and standard deviation. Bold values indicate the best method. Superscripts indicate grouping using Tukey’s HSD test.
@@
None DualLQR 0.96$\pm$0.11a 0.59$\pm$0.039a 1.69$\pm$0.073a
InfLQR 0.91$\pm$0.15ab 0.62$\pm$0.033b 1.77$\pm$0.076c
Single LQR 0.89$\pm$0.23b 0.68$\pm$0.048c 1.73$\pm$0.060b
Low, Orientation DualLQR 0.95$\pm$0.16a 0.61$\pm$0.041a 2.42$\pm$0.119a
InfLQR 0.83$\pm$0.23a 0.63$\pm$0.033a 2.47$\pm$0.114a
Single LQR 0.89$\pm$0.21a 0.77$\pm$0.138b 2.71$\pm$0.208b
Low, Position DualLQR 0.98$\pm$0.10a 0.73$\pm$0.043a 1.69$\pm$0.072a
InfLQR 0.90$\pm$0.15a 0.76$\pm$0.039a 1.78$\pm$0.075b
Single LQR 0.90$\pm$0.23a 0.92$\pm$0.123b 1.74$\pm$0.118ab
Medium, Orientation DualLQR 0.92$\pm$0.15a 0.66$\pm$0.081a 3.41$\pm$0.261a
InfLQR 0.72$\pm$0.20b 0.66$\pm$0.039a 3.39$\pm$0.192a
Single LQR 0.89$\pm$0.20a 0.91$\pm$0.088b 3.85$\pm$0.487b
Medium, Position DualLQR 0.95$\pm$0.16a 0.94$\pm$0.102a 1.69$\pm$0.090a
InfLQR 0.91$\pm$0.19a 1.00$\pm$0.046b 1.79$\pm$0.080b
Single LQR 0.92$\pm$0.20a 1.31$\pm$0.088c 1.73$\pm$0.060a
High, Orientation DualLQR 0.88$\pm$0.13a 0.65$\pm$0.044a 4.44$\pm$0.279a
InfLQR 0.55$\pm$0.25b 0.69$\pm$0.053a 4.28$\pm$0.271a
Single LQR 0.77$\pm$0.24a 1.05$\pm$0.101b 5.28$\pm$0.394b
High, Position DualLQR 0.96$\pm$0.14a 1.20$\pm$0.078a 1.69$\pm$0.094a
InfLQR 0.91$\pm$0.14a 1.23$\pm$0.054a 1.81$\pm$0.078b
Single LQR 0.91$\pm$0.21a 1.75$\pm$0.084b 1.72$\pm$0.064a
5.2 Results of apple grasping experiment
When testing the performance of DualLQR on apple grasping, we found a very high success rate. Only a single grasp attempt out of 110 was unsuccessful, at a control cost of $-0.3$, shown with a cross in Fig. 6. For this control cost, the final approach accuracy was also insufficient, as shown in Fig. 6. From -0.9, the required accuracy was met, except for -2.1. At this control cost, the yaw of the apple was around $0.3$ less than the other settings.
@@ 5.2 Results of apple grasping experiment
In this experiment, the control cost needed to be lower to achieve the required accuracy, as shown in the left plot in Fig. 6. However, similar accuracy was achieved with low control costs. Furthermore, the distance travelled by the manipulator was nearly doubled compared to the simulation, as shown in the middle plot in Fig. 6. Lastly, the right plot of Fig. 6 shows that the time until successful grasping was the lowest at a control cost of $-1.2$. At a control cost of $-1.8$ or lower, there was a lot of variation, while control costs above $-1.2$ were slightly slower.